The angle of incidence of a light ray on a mirror surface is 35 degrees. What is the angle between the incident and reflected rays?




A beam of light is incident on a plane mirror at an angle of 32 degrees relative to the normal line. What is the angle between reflected rays and the surface mirrors?

1
Expert's answer
2022-04-20T13:46:51-0400

1.

"\gamma=2\alpha=70°."

2.

"\theta=90°-\beta=90°-\alpha=58°."